OpenStudy (anonymous):

find all SOLUTIONS of the following equation in the interval [0,2π) tan x- cot x=0

OpenStudy (festinger):

tan x = sin x/cos x cot x = 1/tanx therefore the equation reduces to: \[\frac{sin x}{cos x}=\frac{cos x}{sin x}\] and thus \[cos^2 x - sin^2 x = 0\] using the identity: \[cos^2 x - sin^2 x = cos2x\] It becomes: \[cos2x=0\] Thus, \[2x=\frac{\pi}{2}+n\pi\]